OpenStudy (unknownrandom):

Can somebody show me how to work this out? x^3-2x^2+3x+10=0

OpenStudy (anonymous):

So you need to find the roots?

OpenStudy (unknownrandom):

Yes ma'am @raechelvictoria

OpenStudy (unknownrandom):

So far I found the factors of 10 and x^3. Am I headed in the right direction?

OpenStudy (anonymous):

Yes. Then substitute those values into the equation until you the equation will equal to zerio (remainder theorem)

OpenStudy (unknownrandom):

I couldn't find a number that balanced it out. Maybe I messed up on the math. I will try again.

zepdrix (zepdrix):

Hmm it turns out the real root is x = -1.3387. Are you sure you posted the equation correctly?

OpenStudy (unknownrandom):

I'm dumb I posted the wrong thing. Here is the original equation: q(x)=x^4-4x^3+x^2+16x-20. That other equation was after synth div

OpenStudy (unknownrandom):

When 2 was under the bracket. Where do I go from there?

zepdrix (zepdrix):

Oh so you already found 2 to be a root? That looks correct. Maybe you made a boo boo when you did the division, lemme check.

zepdrix (zepdrix):

Synthetic Division is giving me: x^3 - 2x^2 `- 3x` + 10 = 0 See the minus on the 3x? Maybe you made a small error right there.

OpenStudy (unknownrandom):

Indeed. Thank you for catching that. What next?

zepdrix (zepdrix):

So you took the factors of your constant and divided them by the factors of your leading coefficient. That gave you possible rational roots. You'll need to check them again :c grr Hint hint: Check 2 and -2 again. One of those should work out.

OpenStudy (unknownrandom):

I'm confused what you mean I need to check them again?

OpenStudy (unknownrandom):

I see what you mean @zepdrix -2 did work so now I'm going to use the quad formula. Thanks for all the help!

OpenStudy (unknownrandom):

I ended up with roots of -2,2,2+i, and 2-i

zepdrix (zepdrix):

Ah good job! :)

OpenStudy (unknownrandom):

This is another one I'm going to work on g(x)=x^6-1. So would I do the same steps?

zepdrix (zepdrix):

For g(x) we can write it as the `difference of squares`. x^6 = (x^3)^2 1 = 1^2 x^6 - 1 = (x^3)^2 - 1^2 Do you remember how to factor the difference of squares? They break down into `conjugates`.

OpenStudy (unknownrandom):

No I don't remember how to do that.

zepdrix (zepdrix):

Difference of squares break down into factors like this: a^2 - b^2 = (a-b)(a+b) So in our problem, it would factor down to: (x^3 - 1)(x^3 + 1)

OpenStudy (unknownrandom):

I ended up doing syth division, but that other way is quicker.

zepdrix (zepdrix):

Oh synthetic worked on this one? cool

OpenStudy (unknownrandom):

So here are my final roots 1,-1, -1+-iSqrt(3)/2, and 1+-isqrt(3)/2. Does this look right?

zepdrix (zepdrix):

Ya that's what I'm getting also.

OpenStudy (unknownrandom):

Alright, thanks again for helping me out!

OpenStudy (anonymous):

Refer to the Mathematica calculation.

OpenStudy (unknownrandom):

Thank you both @robtobey @zepdrix

OpenStudy (anonymous):

Your welcome. You might be interested in the following video. http://www.ted.com/talks/conrad_wolfram_teaching_kids_real_math_with_computers.html